THE PARTITION FUNCTION, Z_{H}, SPECIFIC TO THE DIATOMIC MOLECULAR GAS INTERACTION WITH THE LASER RADIATION - THE STUDY OF THE MULTIPHOTONIC EXCITES FOR THE MOLEC-ULAR DISSOCIATIONS

\underline{S. POPESCU}, Research Institute for Electrical Engineering, ICPE - Bistrita, RO-4400 Romania.